OpenStudy (anonymous):

when i need to solve a polynomial equation like : (x^2+7x^2-2x-14)(x^3+x^2+2x+2)=0 do i need to factor both sides then do long division or what.. ???

OpenStudy (anonymous):

i think you can just divide one of the polynomials to both sides so you only deal with one of them

OpenStudy (anonymous):

im not sure i understand

OpenStudy (anonymous):

(x^2+7x^2-2x-14)(x^3+x^2+2x+2)=0 (x^2+7x^2-2x-14)=0*(1/(x^3+x^2+2x+2)) (x^2+7x^2-2x-14)=0

OpenStudy (anonymous):

oh ok so i got rid of one of those (.....) and now i can just find the zeros of this one i have left.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

and use that to do long div. and solve

OpenStudy (anonymous):

understood. thanks

OpenStudy (anonymous):

factor out or complete the square or quadratic formula which ever you like.

OpenStudy (anonymous):

or plug into calculator?

OpenStudy (anonymous):

if your calc has solve for function

OpenStudy (anonymous):

http://www.numberempire.com/equationsolver.php

OpenStudy (anonymous):

i was just gonna plug it in. graph find the zeros. and use that as the divisor. for long div

OpenStudy (anonymous):

you can graph and finds zeros :) yes but you could have done that from the beginning withouth having to have divided by both sides :P

OpenStudy (anonymous):

so when im given two things in parentheses = to zero.. i need to graph both of them or just one

OpenStudy (anonymous):

It is a good idea to do both :)

OpenStudy (anonymous):

First one then the other

OpenStudy (anonymous):

ahh. ok i see, thanks

OpenStudy (anonymous):

so to solve it can i do